6. Suppose a certain contractor wants to maximize profit from building one bridge. The contractor uses both labor and capital and efficient combinations of Labor and capital that are sufficient to make a bridge is given by the function 0.25 L0.5 K0.5. If the prices of labor and capital are birr 5 and birr 10 respectively. Find the least cost combination of L and K and the minimum cost.

1
Expert's answer
2022-02-01T16:01:03-0500

MRTS = "\frac{K}{L}"


"\frac{K}{L} = \frac{w}{r}"


"\frac{K}{L} = \frac{6}{16}"


"\frac{K}{L} = 0.375"


K = 0.375L

The least cost combination of L and K is where: K = 0.375L